OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points) Thanks a lot

OpenStudy (anonymous):

eq of the line having slope m and passing through pt (x1,y1) in point-slope form is given by y-y1=m(x-x1)

OpenStudy (anonymous):

here slope is 4 and x1=3 and y1=-6

OpenStudy (anonymous):

no i still don't understand i have done read the lesson multiple times and i don't understand

OpenStudy (anonymous):

see slope is the tangent of the inclination angle which the given line makes with the +ve direction of x-axis

OpenStudy (anonymous):

so the equation would be y-6=4(x-3)

OpenStudy (anonymous):

for part a) eq f the reqd line is y-(-6)=4(x-3)

OpenStudy (anonymous):

u did it correctly except y+6 instead of y-6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i know slope intercept is y=mx+b

OpenStudy (anonymous):

but how do you get read of the x1 and y1

OpenStudy (anonymous):

rid

OpenStudy (anonymous):

x1,y1 is the cordinate of the point through which the line passes

OpenStudy (anonymous):

right but that is not in y=mx+b

OpenStudy (anonymous):

y=mx+ b is the slope-interept form

OpenStudy (anonymous):

yes but it says make the first equation into slope intercept form

OpenStudy (anonymous):

after simplifying just rearrange the terms keeping y on the left side of equal sign

OpenStudy (anonymous):

y+6=4(x-3) y+6=4x-12 y=4x-18 is that right

OpenStudy (anonymous):

y-(-6)=4(x-3) or y+6= 4x -12 y=4x-12-6 y=4x-18 (point slope form)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okay i think i got in now so for standard form y=4x-18 -4x+y=-18

OpenStudy (anonymous):

also for standard form just rearrange the terms u may write 4x-y-18=0 or 4x-y=18 as per ur choice

OpenStudy (anonymous):

wait for the second one how come your 4x isn't negative